500.0 g of water is warmed to 84.5 degrees Celsius after 77.0 kJ of heat have been added to it. What was the initial temperature of the water?
1
Expert's answer
2017-02-02T13:00:06-0500
According to Q = mC(T2 - T1), T1 = T2 - Q/mC = 47.7 C.